OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x + y + 2z = 1 x + 2y – z = –8 3x – y – 3z = 18

OpenStudy (primeralph):

x = -8 -2y +z

OpenStudy (anonymous):

7y – z = 25 5y – 6z = 42 –5x – 5z = –10 6x – z = 19 7x + 3z = –10 –5z = 19 –5y + 5z = 25 –7y = 42

OpenStudy (anonymous):

those are my options A,B,C,D

OpenStudy (primeralph):

Cool, I'm not done yet.

OpenStudy (primeralph):

@johnweldon1993 Please take over. Got to go.

OpenStudy (anonymous):

thankyou anyway @primeralph :)

OpenStudy (johnweldon1993):

So isolate 'x' in that 2nd equation x + 2y – z = –8 subtract 2y and add z to both sides to isolate x x = -2y + z - 8 Now when you substitute that into the other 2 equations you get 3(-2y + z - 8) + y + 2z = 1 3(-2y + z - 8) – y – 3z = 18 simplified down we get -6y + 3z - 24 + y + 2z = 1 -5y + 5z - 24 = 1 -5y + 5z = 25 And -6y + 3z - 24 - y - 3z = 18 -7y - 24 = 18 -7y = 42 So those 2 again are -5y + 5z = 25 -7y = 42 Look familiar? :)

OpenStudy (anonymous):

Thankyou so much so answer is D :)

OpenStudy (johnweldon1993):

That is correct! :)